Circular No. 6001

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA OPHIUCHI 1994 S. Nakano, Sumoto, Japan, reports the discovery by Akihiko Tago

(Yanahara-cho, Kume-gun, Okayama) of an apparent nova of mag 7.0 on T-Max 400 exposures taken with a 50-mm f/2.8 camera lens on June 1.657 and 2.672 UT; the star was not visible on his patrol films through May 20. Nakano forwards the following position from Y. Kushida, Yatsugatake South Base Observatory: α = 17^{h}35^{m}44^{s}.61, δ = −19^{o}19′34′′.1 (equinox 2000.0); the corresponding 1950.0 position is α = 17^{h}32^{m}47^{s}.56, δ = −19^{o}17′41′′.8. A measurement from A. Sugie, Dynic Astronomical Observatory, gives the end figures as 44^{s}.62 and 33′′.9 (equinox 2000.0). Additional magnitude estimates: May 16.63, [9 (M. Sugano, Kakogawa, Hyogo; Tri-X film); June 1.73, 7.5–8 (Sugano); 2.621, 6.5 (Minoru Yamamoto, Okazaki, Aichi; independent discovery on T-Max film); 3.54, 8.5 (Kushida; visual); 3.557, 7.9 (Kushida; T-Max 400 film); 3.617, 7.8 (S. Takahashi, Dynic Astronomical Observatory; visual).

SUPERNOVA 1994Q IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur (OCA), reports his discovery

of a supernova (mag R ∼ 16.5) at α = 16^{h}48^{m}11^{s}.64, δ = +40^{o}31′01′′.1 (equinox 1950.0), which is 0′′.4 west and 3′′.5 south of the center of the host galaxy. SN 1994Q was found on two plates taken by D. Albanese on June 2.92 and 3.06 UT with the 0.9-m OCA Schmidt telescope; nothing is visible at this position on a similar plate taken one month ago. A nearby star (R ∼ 15) has end figures 05^{s}.97, 30′50′′.5. A. V. Filippenko, T. Matheson, and A. J. Barth, University of California at Berkeley, report that inspection of uncalibrated CCD spectra (range 310–800 nm, resolution 0.6 nm) obtained on June 3 with the Shane 3-m reflector at Lick Observatory suggests that the object is a supernova of type Ia, probably ∼ 1 week past maximum brightness.

COMET TAKAMIZAWA (1994i) Total visual magnitude estimates (cf. IAUC 5987): May 13.14 UT, 9.3

(J. E. Bortle, Stormville, NY, 20×80 binoculars); 15.99, 9.7 (L. Szentaskó, Veresegyház, Hungary, 0.33-m reflector); June 2.28, 9.8 (C. E. Spratt, Victoria, BC, 0.20-m reflector).

1994 June 3 Daniel W. E. Green